# MathCAMPS: Fine-grained Synthesis of Mathematical Problems From Human Curricula

### v1

@@ Appendix A Common Core Standards in MathCAMPS
Table 4: CC Standards for Grade K
@@
Table 5: CC Standards for Grade 1
@@
Table 6: CC Standards for Grade 2
@@
Table 7: CC Standards for Grade 3
@@
4.NF.A.2 Compare two fractions with different numerators and different denominators, e.g., by creating common denominators or numerators, or by comparing to a benchmark fraction such as 1/2. Recognize that comparisons are valid only when the two fractions refer to the same whole. Record the results of comparisons with symbols >, =, or <, and justify the conclusions, e.g., by using a visual fraction model.
@@
Table 8: CC Standards for Grade 4
@@
Table 9: CC Standards for Grade 5
@@
Table 10: CC Standards for Grade 6
@@
Table 11: CC Standards for Grade 7
@@
Table 12: CC Standards for Grade 8
Appendix B Familiarity bias
GPT-4o 0.910 0.954
Claude 3 Opus 0.887 0.909
Appendix C Data generation pipeline details
C.1 Grammar
@@
C.2 Answer Grading During Evaluation
@@ C.3 Cost estimate
Appendix D Correlation between MathCAMPS and GSM8k
@@
Figure 3: Relation between accuracy on GSM8k and on MathCAMPS.
Appendix E Largest Model Rank Changes When Focusing on One CC Standard (Complete Table)
Table 14 shows the full table from which Table 2 was extracted.
@@ Appendix E Largest Model Rank Changes When Focusing on One CC Standard (Complete Table)
Claude-3 Opus 2.MD.B.5 - Add/sub within 100 ($2^{nd}$ $13^{th}$)
Gemini-1.5 Pro K.OA.A.4 - Adding to equal 10 ($3^{rd}$ $19^{th}$)
Gemini-1.5 Flash 4.OA.B.4 - Factor pairs within 100 ($4^{th}$ $20^{th}$)
GPT-3.5 Turbo 6.EE.A.1 - Evaluate exponents ($5^{th}$ $21^{th}$)
Claude-3 Sonnet 2.NBT.B.5 - Add/sub within 100 ($6^{th}$ $12^{th}$)
Claude-3 Haiku 3.OA.A.4 - Determine unknowns in mul/div probs ($9^{th}$ $1^{st}$)
Llama 3 70B K.OA.A.4 - Adding to equal 10 ($7^{th}$ $17^{th}$)
Mixtral 8x22B 8.EE.C.8 - Solve two-variable systems ($8^{th}$ $21^{th}$)
DeepSeek 67B K.NBT.A.1 - Decompose into 10s ($10^{th}$ $1^{st}$)
Llama 3 8B 4.NBT.B.4 - Add/sub multi-digit nums ($11^{th}$ $21^{th}$)
Mixtral 8x7B 6.EE.A.1 - Evaluate exponents ($12^{th}$ $20^{th}$)
Llemma 34B K.OA.A.4 - Adding to equal 10 ($13^{th}$ $1^{st}$)
Mistral 7B 1.OA.A.1 - Add/sub within 20 ($14^{th}$ $21^{th}$)
DeepSeek Coder 33B 6.EE.A.1 - Evaluate exponents ($15^{th}$ $3^{rd}$)
CodeLlama 34B 5.NF.A.1 - Add/sub fractions ($16^{th}$ $22^{th}$)
phi-2 K.OA.A.4 - Adding to equal 10 ($17^{th}$ $4^{th}$)
Llemma 7B 6.EE.A.1 - Evaluate exponents ($18^{th}$ $5^{th}$)
Gemma 7B K.OA.A.5 - Add/sub within 5 ($19^{th}$ $6^{th}$)
CodeLlama 7B 8.EE.C.8 - Solve two-variable systems ($21^{th}$ $15^{th}$)
Gemma 2B 8.EE.C.8 - Solve two-variable systems ($22^{th}$ $11^{th}$)
Table 14: Largest changes in a model’s ranking when comparing its performance in a particular CC standard, in contrast to only overall performance. This is a complete version of Table 2, which only showed some models for brevity.
Appendix F Followup Analysis
Table 15 lists model accuracies when only looking at the main problems (Main Acc.), their accuracies when only looking at the incremental followups (IFUP Acc.), their accuracies when only looking at the counterfactual followups (CFUP Acc.), and finally, the total number of followups seen by each model. The total number of followups a model sees relies on whether or not they get the main question for that followup correct. If a model does not correctly solve the main question, it is not prompted with follow-ups. Note that each followup serves as a followup to the main question, as opposed to a followup to each other.
Vendor Model Main Acc. IFUP Acc. CFUP Acc. Total FUPs seen
Anthropic Claude-3 Opus 0.89 0.90 0.88 4142
Anthropic Claude-3 Sonnet 0.86 0.86 0.87 3964
Anthropic Claude-3 Haiku 0.84 0.88 0.87 3819
DeepSeek DeepSeek Coder 33B 0.65 0.79 0.85 1022
DeepSeek DeepSeek 67B 0.80 0.87 0.88 3286
EleutherAI LLemma 7B 0.62 0.68 0.80 2890
Google Gemini-1.5 Pro 0.89 0.91 0.89 4140
Google Gemini-1.5 Flash 0.87 0.89 0.87 4083
Google Gemma 2B 0.51 0.29 0.54 2044
Google Gemma 7B 0.62 0.55 0.60 2786
Meta Llama 3 8B 0.77 0.84 0.80 3476
Meta Llama 3 70B 0.85 0.87 0.84 3939
Meta CodeLlama 7B 0.52 0.69 0.86 617
Meta CodeLlama 13B 0.58 0.75 0.80 2451
Meta CodeLlama 34B 0.64 0.82 0.88 844
Microsoft phi-2 0.63 0.48 0.78 2873
Mistral Mistral 7B 0.68 0.72 0.80 3090
Mistral Mixtral 8x7B 0.76 0.80 0.82 3439
Mistral Mixtral 8x22B 0.84 0.86 0.83 3948
OpenAI GPT-4o 0.92 0.92 0.90 4358
OpenAI GPT-3.5 Turbo 0.87 0.85 0.86 4063
Table 15: Accuracy of each model on incremental follow-up questions (IFUP) as well as on counterfactual follow-ups (CFUP). Note that these accuracies are not directly comparable, since models are only evaluated on follow-ups to problems that they respond correctly to; thus, each accuracy shown here is over a different subset of follow-up problems in MathCAMPS.

### v2

@@ Appendix C Common Core Standards in MathCAMPS
Table 1: $C$ Standards for Grade K
@@
Table 2: $C$ Standards for Grade 1
@@
Table 3: $C$ Standards for Grade 2
@@
Table 4: $C$ Standards for Grade 3
@@
4.NF.A.2 Compare two fractions with different numerators and different denominators, e.g., by creating common denominators or numerators, or by comparing to a benchmark fraction such as 1/2. Recognize that comparisons are valid only when the two fractions refer to the same whole. Record the results of comparisons with symbols ¿, =, or ¡, and justify the conclusions, e.g., by using a visual fraction model.
@@
Table 5: $C$ Standards for Grade 4
@@
Table 6: $C$ Standards for Grade 5
@@
Table 7: $C$ Standards for Grade 6
@@
Table 8: $C$ Standards for Grade 7
@@
Table 9: $C$ Standards for Grade 8
Appendix D Data generation pipeline details
D.1 Grammar
@@
D.2 Answer Grading During Evaluation
@@ D.3 Cost estimate
Appendix E Correlation between MathCAMPS and GSM8k
@@
Figure 18: Correlation between accuracy on GSM8k and on MathCAMPS ($\rho=0.91$).
Appendix F Cycle consistency efficacy and failure cases
This cycle consistency test significantly improves the reliability of our pipeline. We manually evaluated 245 random problems generated by sampling a symbolic structure and then a word problem from GPT-4. Out of those, we identified 30 word problems (12.2%) that were not faithful to the original symbolic structure — for those, the answer that we compute to the symbolic problem does not match our manual solution to the word problem. Cycle consistency discarded 25 of those (and 7 problems that were indeed faithful). Out of the remaining 215 problems, 210 (97.7%) were judged as faithful in our manual check. The results are shown in Table 10.
Faithful problem Unfaithful problem
Cycle-consistent 208 5
Not cycle-consistent 7 25
Table 10: Efficacy of Cycle Consistency
F.1 Unfaithful but cycle-consistent problems
We noted that of the five problems that were unfaithful but cycle-consistent, two had ambiguous language that led to the generation of a cycle-consistent symbolic structure. The three other problems were all from the same CC standard (4.NBT.B.6, a skill regarding the calculation of quotients and remainders).
F.2 Unfaithful and non-cycle consistent problems
Below, we provide threeexamples of unfaithful problem generations that were unfaithful and correctly discarded by the cycle-consistency process.
Original symbolic structure: [[var f = (5 - 4)]][[var t = (4 - f)]][[question d = [’t’]]] theme: Rocket
Generated word problem: In the space exploration research center, Diego has 5 rockets. He launched 1 rocket for testing purposes. After this, he orders 4 new rockets, but doesn’t count the one that he just launched. How many rockets does Diego have now?
New symbolic structure: [[var initial_rockets = 5]][[var launched_rockets = 1]][[var ordered_rockets = 4]][[var total_rockets = (initial_rockets - launched_rockets + ordered_rockets)]][[question result = [’total_rockets’]]]
Original symbolic structure: [[var v = (79.24 * 37.6)]][[question s = [’v’]]] theme: Treasure chest
Generated word problem: A pirate finds a treasure chest full of golden coins. Each golden coin weighs 79.24 grams. If the total weight of the coins is 37.6 kilograms, how many golden coins are there in the treasure chest?
New symbolic structure: [[var weightInGrams = (37.6 * 1000)]][[var n = (weightInGrams / 79.24)]][[question numCoins = [’n’]]]
Original symbolic structure: [[var j = (4450 / 89)]][[var c = ((69 * j) + j)]][[question b = [’c’]]] theme: Sandwich
Generated word problem: The Sandwich Factory makes sandwiches at a rate of 4450 sandwiches every 89 days. If the factory not only produces 69 sandwiches each day, but also an additional batch of the exact same rate each day, how many sandwiches are produced in total each day?
New symbolic structure: [[var a = 4450]][[var b = 89]][[var c = 69]][[var d = a / b]][[var e = c + d]][[question f = [’e’]]]
F.3 Faithful but non-cycle-consistent problems
The 7 faithful problems were discarded were because one of two issues. 6 our of 7 problems were discarded because their back-translated symbolic structure had a small structural error (e.g. and extra square bracket at the end of a variable declaration). The 7th problem had a genuine error in its back translation, which caused the new symbolic structure to have a different final answer than the original symbolic structure, causing us to discard the problem.
Appendix G Familiarity bias
GPT-4o 0.910 0.954
Claude 3 Opus 0.887 0.909
Appendix H Performance of Families of Models on MathCAMPS
Vendor Model All K 1 2 3 4 5 6 7 8
OpenAI GPT-4o 0.92 0.98 0.98 0.98 0.98 0.92 0.88 0.95 0.89 0.64
Anthropic Claude-3 Opus 0.89 0.97 0.99 0.96 0.98 0.89 0.83 0.96 0.73 0.56
Google Gemini-1.5 Pro 0.89 0.95 0.98 0.97 0.97 0.89 0.83 0.93 0.78 0.54
Google Gemini-1.5 Flash 0.87 0.98 0.98 0.97 0.98 0.80 0.80 0.90 0.84 0.56
OpenAI GPT-3.5 Turbo 0.87 0.96 0.98 0.98 0.97 0.86 0.77 0.90 0.77 0.56
Anthropic Claude-3 Sonnet 0.86 0.96 0.98 0.97 0.98 0.88 0.74 0.94 0.66 0.49
Anthropic Claude-3 Haiku 0.84 0.97 0.98 0.97 0.98 0.87 0.69 0.92 0.59 0.51
Qwen Qwen2-Math 72B 0.89 0.98 0.99 0.98 0.97 0.90 0.80 0.91 0.77 0.59
Meta Llama 3 70B 0.85 0.96 0.97 0.97 0.97 0.85 0.71 0.87 0.73 0.50
Mistral Mixtral 8x22B 0.84 0.96 0.99 0.98 0.96 0.79 0.69 0.88 0.73 0.61
Qwen Qwen2-Math 7B 0.83 0.96 0.99 0.97 0.93 0.85 0.66 0.91 0.58 0.62
DeepSeek DeepSeek 67B 0.80 0.95 0.99 0.96 0.93 0.82 0.60 0.84 0.61 0.47
DeepSeek DeepSeek Math 7B Base 0.78 0.94 0.97 0.93 0.89 0.75 0.63 0.86 0.53 0.55
Numina NuminaMath 7B TIR 0.78 0.89 0.97 0.95 0.90 0.72 0.63 0.84 0.59 0.53
Meta Llama 3 8B 0.77 0.94 0.97 0.96 0.94 0.78 0.55 0.79 0.53 0.43
Mistral Mixtral 8x7B 0.76 0.94 0.96 0.93 0.91 0.75 0.52 0.80 0.53 0.45
InternLM InternLM-Math Base 20B 0.74 0.95 0.96 0.95 0.86 0.68 0.55 0.79 0.52 0.47
EleutherAI Llemma 34B 0.71 0.95 0.96 0.93 0.87 0.61 0.47 0.77 0.46 0.44
Mistral Mistral 7B 0.68 0.89 0.94 0.91 0.84 0.61 0.42 0.66 0.45 0.42
DeepSeek DeepSeek Coder 33B 0.65 0.88 0.93 0.92 0.83 0.54 0.36 0.66 0.44 0.38
Meta CodeLlama 34B 0.64 0.90 0.94 0.92 0.85 0.51 0.38 0.70 0.37 0.30
Microsoft phi-2 0.63 0.95 0.96 0.89 0.78 0.46 0.38 0.61 0.37 0.41
EleutherAI Llemma 7B 0.62 0.78 0.90 0.85 0.79 0.48 0.41 0.67 0.41 0.36
Google Gemma 7B 0.62 0.83 0.92 0.90 0.82 0.47 0.36 0.65 0.36 0.30
Meta CodeLlama 13B 0.58 0.87 0.92 0.87 0.75 0.41 0.30 0.61 0.32 0.34
InternLM InternLM-Math Base 7B 0.58 0.71 0.73 0.73 0.72 0.54 0.38 0.61 0.37 0.39
Meta CodeLlama 7B 0.52 0.85 0.92 0.84 0.69 0.37 0.25 0.57 0.25 0.16
Google Gemma 2B 0.51 0.66 0.76 0.74 0.67 0.42 0.28 0.55 0.30 0.27
- Avg. Performance 0.75 0.91 0.95 0.92 0.88 0.70 0.57 0.79 0.56 0.46
Table 12 shows both aggregate accuracy on MathCAMPS, as well as accuracy across standards partitioned by grade, whereas Figure 18 compares the aggregate accuracies on MathCAMPS and GSM8K. Closed-weights models are shown above the line, with open-weights models below. GPT-4o ranks at the top in overall accuracy. Since we used GPT-4 to generate the problems, we must rule out familiarity bias (Stureborg et al., 2024) in this result. We thus generated a 10%-scale dataset with the same pipeline but using Claude-3 Opus. We found that GPT-4o still outperforms Claude-3 Opus on this dataset (see Appendix G), suggesting that its advantage on MathCAMPS was not due to a familiarity bias.
We make the following observations:
@@ Appendix H Performance of Families of Models on MathCAMPS
Claude-3 Opus 2.MD.B.5 - Add/sub within 100 ($2^{nd}$ $18^{th}$)
Gemini-1.5 Pro K.OA.A.4 - Adding to equal 10 ($4^{th}$ $23^{th}$)
Claude-3 Haiku 6.EE.A.1 - Evaluate exponents ($10^{th}$ $20^{th}$)
Llama 3 70B 3.OA.A.3 - Mul/div within 100 ($8^{th}$ $21^{th}$)
Mixtral 8x22B 8.EE.C.8 - Solve two-variable systems ($9^{th}$ $21^{th}$)
Qwen2-Math 7B 8.EE.C.8 - Solve two-variable systems ($11^{th}$ $25^{th}$)
DeepSeek 67B K.NBT.A.1 - Decompose into 10s ($12^{th}$ $1^{st}$)
Llama 3 8B K.OA.A.4 - Adding to equal 10 ($15^{th}$ $3^{rd}$)
Mixtral 8x7B 6.EE.A.1 - Evaluate exponents ($16^{th}$ $26^{th}$)
InternLM-Math Base 20B 2.NBT.B.5 - Add/sub within 100 ($17^{th}$ $2^{nd}$)
Llemma 34B 3.OA.A.3 - Mul/div within 100 ($18^{th}$ $1^{st}$)
Mistral 7B 1.OA.A.1 - Add/sub within 20 ($19^{th}$ $26^{th}$)
DeepSeek Coder 33B 6.EE.A.1 - Evaluate exponents ($20^{th}$ $3^{rd}$)
phi-2 K.OA.A.4 - Adding to equal 10 ($22^{th}$ $4^{th}$)
Llemma 7B 6.EE.A.1 - Evaluate exponents ($23^{th}$ $5^{th}$)
Gemma 7B K.OA.A.5 - Add/sub within 5 ($24^{th}$ $6^{th}$)
InternLM-Math Base 7B 4.OA.B.4 - Factor pairs within 100 ($26^{th}$ $15^{th}$)
CodeLlama 7B 8.EE.C.8 - Solve two-variable systems ($27^{th}$ $15^{th}$)
Gemma 2B 8.EE.C.8 - Solve two-variable systems ($28^{th}$ $11^{th}$)
When considering overall performance, the trends in GSM8k hold on the novel problems from MathCAMPS, which cover overlapping topics (Pearson correlation of 0.865, $p<10^{-5}$; we show this correlation in Figure 18). This correlation corroborates the progress that public benchmarks have witnessed, suggesting that data contamination does not play a major role in explaining observed improvements in recent LLMs. We note that prior work attempting to replicate the distribution of GSM8k, such as the independent effort to collect GSM1k (Zhang et al., 2024), has observed a smaller correlation, including substantial drops in performance for some models. This is entirely compatible with our findings here, due to the difficulty of exactly replicating the distribution over skills in any given human-created benchmark. As the sharp differences in Table 13 indicate, an (unintended) shift in this distibution can drastically — and unevenly — affect accuracy, even if no data contamination occurs. These shifts are easily avoided in an automated pipeline as in MathCAMPS, allowing us to draw new problems from the exact same distribution in the future.
H.1 Standard-specific analysis
Despite decently high performance across the board, GPT-4o’s performance fell at or below $90\%$ on the following skills: 4.MD.A.2-fraction, 4.OA.A.3, 5.NF.A.1, 7.NS.A.3-fraction, and 8.EE.C.8. At their core, all these abilities require fraction addition or subtraction, a skill we noted that GPT-4o struggles with. Specifically, the model starts approximating fractions using decimals, and the error introduced by this compounds throughout the problem, resulting in an incorrect final answer. Surprisingly, GPT-4o achieves an $86\%$ on 5.NF.B.4, which requires fraction multiplication, indicating that it is likely the multi-step process of finding common denominators in adding/subtracting fractions that challenges GPT-4o. Additionally, GPT-4o achieves performances above $90\%$ on 4.MD.A.2-decimal and 7.NS.A.3-decimal, which are the CC standards equivalent to 4.MD.A.2-fraction and 7.NS.A.3-fraction, using decimals instead of fractions in the problems. This trend isn’t isolated to the GPT models, though, as most models tended to struggle more with standards involving fractions.
Work from Lucy et al. (2024) showed that over $50\%$ of problems from GSM8K originated from three CC standards, namely, 4.OA.A.3 ($20.73\%$), 2.OA.A.1 ($16.58\%$), and 3.OA.D.8 ($15.75\%$). These standards ask students to solve multistep word problems involving the four operations, use addition and subtraction to solve two-step word problems within 100, and solve two-step word problems using the four operations, respectively. While most models we experimented with performed relatively well on 2.OA.A.1 and 3.OA.D.8, CC standard 4.OA.A.3 did prove to be challenging, with the most performant model, Qwen2-Math 72B, achieving an $86\%$ on the standard.
Out of the 49 total skills we evaluated (44 standards, some of which we split into sub-standards), 19 skills had an absolute winner: a model which outperforms all other models on that skill. The distribution of these skills is given in Table 14. This analysis shows that even generally weaker models, such as GPT-3.5 Turbo, have particular skills that they excel on. This fact is hidden when looking at aggregate accuracies, but is revealed in our finer-grained analysis.
Model Standards Won
GPT-4o 4.NBT.B.6, 7.NS.A.2, 8.EE.C.7, 7.NS.A.1-fraction, 5.NF.A.1, 7.NS.A.3-fraction
Qwen2-Math 72B 1.OA.A.1, 3.OA.D.8, 5.NF.B.4, 4.OA.A.3, 4.MD.A.2-fraction
GPT-3.5 Turbo 2.NBT.B.6, 5.OA.A.1, 8.EE.C.8
Claude-3 Opus 6.NS.B.2, 5.NBT.B.7
Gemini-1.5 Flash 7.NS.A.3-decimal, 5.NF.A.2
Claude-3 Sonnet 3.MD.D.8-polygon
Table 14: Standards with strict winners, i.e., models who strictly outperform all other models on that standard.
H.2 Follow-up tasks
Model Acc. w.f. Largest accuracy drop w/ follow-ups
GPT-4o 0.82 5.NF.A.1 - Add/sub fractions 0.86 0.58)
Claude-3 Opus 0.76 7.NS.A.1-fraction - Add/sub with fractions 0.54 0.23)
Gemini-1.5 Pro 0.77 5.OA.A.1 - Evaluating with parentheses 0.95 0.69)
Claude-3 Haiku 0.70 7.NS.A.2 - Mult/div with fractions 0.55 0.26)
Qwen2-Math 72B 0.78 5.NF.A.1 - Add/sub fractions 0.49 0.23)
Llama 3 70B 0.69 4.NF.A.2 - Compare two fractions 0.99 0.66)
Mixtral 8x22B 0.69 7.NS.A.1-fraction - Add/sub with fractions 0.69 0.17)
Qwen2-Math 7B 0.71 5.NF.A.2 - Add/sub fraction word problems 0.41 0.17)
DeepSeek Math 7B Base 0.65 5.NF.B.4 - Mult fractions 0.81 0.57)
NuminaMath 7B TIR 0.62 5.NF.A.2 - Add/sub fraction word problems 0.44 0.18)
Llama 3 8B 0.58 4.NF.A.2 - Compare two fractions 0.90 0.52)
Mixtral 8x7B 0.58 7.NS.A.2 - Mult/div with fractions 0.60 0.28)
Llemma 34B 0.55 5.NF.B.4 - Mult fractions 0.68 0.31)
Mistral 7B 0.48 7.NS.A.1-decimal - Add/sub with decimals 0.91 0.50)
DeepSeek Coder 33B 0.60 3.OA.A.3 - Mul/div within 100 0.95 0.81)
phi-2 0.39 3.NBT.A.2 - Add/sub within 1000 0.71 0.23)
Llemma 7B 0.42 5.NF.B.4 - Mult fractions 0.58 0.21)
Gemma 7B 0.33 7.NS.A.1-decimal - Add/sub with decimals 0.91 0.32)
InternLM-Math Base 7B 0.42 7.NS.A.1-decimal - Add/sub with decimals 0.82 0.47)
CodeLlama 7B 0.49 2.NBT.B.7 - Add/sub within 100 0.80 0.67)
Gemma 2B 0.24 3.NBT.A.2 - Add/sub within 1000 0.93 0.26)
Table 15: Model performance on our mathematical dialogue task, where the model must answer follow-up questions besides the initial problem. The second column, Accuracy with follow-ups, shows overall success rate across standards that contain follow-up questions, considering a model successful only when it answers a problem and its follow-up questions correctly. The third and fourth columns show the hardest standard for each model when it comes to follow-up questions, showing a standard’s code and abbreviated description, the model’s accuracy ignoring follow-ups, and after follow-ups. We show selected rows here, the complete table can be found in the Appendix.